OpenStudy (anonymous):

solve the equation (3tan^2x-1)(tan^2x-3)=0

OpenStudy (anonymous):

Put separately equal to 0.. \[3\tan^2x-1 = 0\] \[\tan^2x-3 = 0\]

OpenStudy (anonymous):

Now find tanx by taking square root..

OpenStudy (campbell_st):

you need to solve \[3\tan^2(x) -1 = 0\] and \[\tan^2(x) - 3 = 0\] looking at the 1st equation \[\tan^2(x) = \frac{1}{3}\] take the square root of both sides the 2nd equation \[\tan^2(x) = 3\] take the square root of both sides...

OpenStudy (campbell_st):

this will give 4 solutions... 2 for each equation.